# Bias-Corrected Data Synthesis for Imbalanced Learning

### v1

@@ 3.3 Average Treatment Effect Estimation
Then the augmented inverse propensity weighting estimator of ATE given $\hat{\mu}_{1}$, $\hat{\mu}_{0}$ and $\hat{e}$ is provided by
@@
Under Assumption 4, denote $r_{a}=(\mathbb{E}|\hat{\mu}_{a}(\bm{X})-\mu_{a}^{*}(\bm{X})|^{2})^{1/2}$ for $a\in\{0,1\}$ and $r_{e}=(\mathbb{E}|\hat{e}(\bm{X})-e^{*}(\bm{X})|^{2})^{1/2}$. There exist constants $C_{0},C_{1}>0$ depending only on $\eta$ and $M$, such that with probability at least $1-\alpha$,
$|\hat{\tau}^{\rm AIPW}-\tau|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}+\frac{C_{1}}{\eta}(r_{1}+r_{0})r_{e}+C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}\big((r_{1}+r_{0})r_{e}+r_{1}r_{0}\big).$
Corollary 1.
Under Assumptions 3 and 4, suppose the propensity score estimation is obtained from the bias-corrected coefficient estimator $\hat{e}^{\rm bc}(\bm{x})=\sigma(\bm{x}^{\top}\hat{\bm{\beta}}^{\rm bc})$, then the corresponding ATE estimator satisfies
$|\hat{\tau}^{\rm AIPW,bc}-\tau|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}+\frac{C_{1}R}{\eta}(r_{1}+r_{0})r_{\beta}+C_{0}R\sqrt{\frac{\log(4/\alpha)}{2n}}\big((r_{1}+r_{0})r_{\beta}+r_{1}r_{0}\big),$
where $r_{\beta}$ represents the error bound for $\hat{\bm{\beta}}^{\rm bc}$:

### v2

@@ 3.3 Average Treatment Effect Estimation
Then the augmented inverse propensity weighting estimator of ATE given $\widehat{\mu}_{1}$, $\widehat{\mu}_{0}$ and $\widehat{e}$ is provided by
@@
Under Assumption 3.7, denote $r_{a}=(\mathbb{E}|\widehat{\mu}_{a}(\bm{X})-\mu_{a}^{*}(\bm{X})|^{2})^{1/2}$ for $a\in\{0,1\}$ and $r_{e}=(\mathbb{E}|\widehat{e}(\bm{X})-e^{*}(\bm{X})|^{2})^{1/2}$. There exist constants $C_{0},C_{1}>0$ depending only on $\eta$ and $M$, such that with probability at least $1-\alpha$,
$|\widehat{\tau}^{\rm AIPW}-\tau|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}+\frac{C_{1}}{\eta}(r_{1}+r_{0})r_{e}+C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}\big((r_{1}+r_{0})r_{e}+r_{1}r_{0}\big).$
Corollary 3.1.
Under Assumptions 3.6 and 3.7, suppose the propensity score estimation is obtained from the bias-corrected coefficient estimator $\widehat{e}^{\rm bc}(\bm{x})=\sigma(\bm{x}^{\top}\widehat{\bm{\beta}}^{\rm bc})$, then the corresponding ATE estimator satisfies
$|\widehat{\tau}^{\rm AIPW,bc}-\tau|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}+\frac{C_{1}R}{\eta}(r_{1}+r_{0})r_{\beta}+C_{0}R\sqrt{\frac{\log(4/\alpha)}{2n}}\big((r_{1}+r_{0})r_{\beta}+r_{1}r_{0}\big),$
where $r_{\beta}$ represents the error bound for $\widehat{\bm{\beta}}^{\rm bc}$: